If (v-p)^2 = 90 and vp=13, find v^2 + p^2
The answer is 116 but i don't understand how to solve for it.

Answer by nerdybill(7384)

If (v-p)^2 = 90 and vp=13, find v^2 + p^2
.
Starting with:
(v-p)^2 = 90
(v-p)(v-p) = 90
apply FOIL to the left side:
v^2 -vp -vp + p^2 = 90
v^2 -2vp + p^2 = 90
Subsitute in the value for vp:
v^2 -2(13) + p^2 = 90
v^2 -26 + p^2 = 90
v^2 + p^2 = 116